# Fast Multiplication and the PLWE–RLWE Equivalence for an Infinite Family of Cyclotomic Subextensions

### v1

@@ 4.2 Case $n=2^{r}3^{s}$
Both the $\mathsf{DCT}$ and the $\mathsf{IDCT}$ of size $N$ can be computed in $\mathcal{O}(N\log N)$ [bi1999fast, kok1997fast], even for the composite dimension $N=2^{r}3^{s-1}$. This means that we can compute $\mathbf{c}$ and hence $r(x)=p(x)q(x)$ in $\mathcal{O}(N\log N)$.
@@
$\mathbf{d}=\begin{bmatrix}c_{0}+c_{m}\\ c_{1}+c_{m+1}-c_{2m-1}\\ \vdots\\ c_{i}+c_{m+i}-c_{2m-i}\\ \vdots\\ c_{m-1}+c_{m+1}\end{bmatrix}.$
@@ 4.3 DCT modulo a prime number
For $n=2^{r}$ or $n=2^{r}3^{s}$, it is enough to choose a prime $q$ satisfying $q\equiv 1\mod n$, because this implies the existence of a primitive $n$-th root of unity, $\zeta_{n}$, in the finite field $\mathbb{F}_{q}=\mathbb{Z}/q\mathbb{Z}$. In both cases, for all indices $i$ and $j$, we use
$2\cos\left(\frac{2\pi(2j+1)i}{4N}\right)=\zeta_{n}^{(2j+1)i}+\zeta_{n}^{-(2j+1)i}\in\mathbb{F}_{q}.$
@@ 5 PLWE in $\mathbb{Z}[x]/(\Psi_{n}(x))$ via fast base change
Before we can prove Lemma 5.1, we need a few more results regarding the complexity of polynomial evaluation on the Chebyshev nodes $x_{j}$ (5) and Chebyshev interpolation. In [pan1998new, gerasoulis1988fast], the authors describe an algorithm to evaluate a polynomial $p(x)$ of degree at most $N-1$ at the Chebyshev nodes $x_{j}$ with complexity $\mathcal{O}(N\log N)$. In [pan1998new], this result is stated as the following theorem.
@@ 6 Small roots of cyclotomic polynomials modulo a prime
Let $\mathcal{O}_{f}=\mathbb{Z}[x]/(f(x))$ and $R_{q}=\mathcal{O}_{f}/q\mathcal{O}_{f}$ as before. The root-based attacks from [ELOS:2016:RCN] show that if the polynomial $f(x)$ has a small root modulo the reducing prime $q$, then there exists a polynomial time evaluation attack against the decisional PLWE problem. In this section, we compare heuristically cyclotomic fields and their maximal real subfields for attacks against roots of small residue. We show that between the two families of minimal polynomials, $\Phi_{p}(x)$ and $\Psi_{4p}(x)$ of the same degree, the maximal real extensions $\Psi_{4p}(x)$ tend to have less small roots $x\in S=\{\pm 2,\pm 3,\pm 4,\pm 8\}$ when $p$ is large enough. On the other hand, for cyclotomic polynomials of small degree $p<200$ and reducing primes $q<5\cdot 10^{10}$, we did not encounter any small roots in $S$.
The roots in the set $S$ are small powers of two. We also included another small prime 3 because our preliminary computations suggested it might occur as root relatively often. This choice allowed us to study the frequency of the roots in $S$ to see if it varies a lot between the elements of $S$.
@@
  • $q=3329$ and $n=256$ for ML-KEM, formerly known as Kyber [FIPS:2024:203].
  • $q=8380417$ and $n=256$ for ML-DSA, formerly known as Dilithium [FIPS:2024:204].
  • $q=12289$ and $n=512$ in security level 1 and $n=1024$ in security level of Falcon [Falcon].
@@
In practice, for PLWE instances the sparsity of the vulnerable polynomials allows us to pre-emptively test for small zeros for different prime moduli $q$ and choose a safe one. In this regard, the weak polynomials we present are only of theoretical interest in the PQC realm. However, it is worth mentioning that in [BDM:2024:GARB], the authors attack the PLWE problem by using as the modulus the NTRU polynomials with parameters:
@@
The situation is slightly different in the homomorphic encryption setting, where much larger values of both the degree $n$ and the prime modulus $q$ can occur. For instance, within HELib [gouert], a large initial ciphertext modulus of 780 bits is used, which needs a cyclotomic degree of 55831. Thus, the ring is of dimension $\phi(55831)=54000$ achieving 129 bits of security. More specifically, in [albertoreciente, Table 4.2] we can observe that $n$ runs in powers of 2 from 1024 to 131072, and $q$, depending on $n$, runs between 29 to 1920 bits. These parameters refer to power-of-two cyclotomic polynomials.
@@
For primes from $p<1500$, we computed the evaluations of $\Phi_{p}(x)$ and $\Psi_{4p}(x)$ at the elements of the set $S$ and then reduced the evaluations with a varying prime modulus $q$ running as high as $5\cdot 10^{10}$. Note that thank to the result from [BL:2022:RPE, Blanco:2021:VMTR], we could exclude the possibility of $\alpha=\pm 2$ being a root of $\Psi_{4p}(x)$ modulo any prime $q$. Furthermore, the minimal polynomials $\Psi_{4p}(x)$ are even [loper2016resultants] so $\alpha\in S$ is root if and only if $-\alpha\in S$ is a root. Thus, in our numerical search the number of roots at $\alpha$ and $-\alpha$ are the same for $\Psi_{4p}(x)$. Finally, note that for a fixed $p$, the degrees of the two polynomials are equal
@@
The counts for the number of zeros we found for the evaluation points in $S$ is given in Table 1. We did not observe a significant difference between the number of zeros at the small powers of 2 or at the small prime $3$. In Table 1 the number of zeros of $\Psi_{4p}(x)$ at $\alpha$ and $-\alpha$ are equal due to the even parity of polynomials.

### v2

@@ 4.2 Case $n=2^{r}3^{s}$
Both the $\mathsf{DCT}$ and the $\mathsf{IDCT}$ of size $N$ can be computed in $\mathcal{O}(N\log N)$ [35, 27], even for the composite dimension $N=2^{r}3^{s-1}$. This means that we can compute $\mathbf{c}$ and hence $r(x)=p(x)q(x)$ in $\mathcal{O}(N\log N)$.
@@
$\mathbf{d}=\begin{bmatrix}c_{0}+c_{m}\\ c_{1}+c_{m+1}-c_{2m-1}\\ \vdots\\ c_{i}+c_{m+i}-c_{2m-i}\\ \vdots\\ c_{m-1}-c_{m+1}\end{bmatrix}.$
@@ 4.3 DCT modulo a prime number
For $n=2^{r}$, let $M=n$. Then it is enough to choose a prime $q$ satisfying $q\equiv 1\mod M$, because this implies the existence of a primitive $M$-th root of unity, $\zeta_{M}$, in the finite field $\mathbb{F}_{q}=\mathbb{Z}/q\mathbb{Z}$. For $n=2^{r}3^{s}$, we choose a prime $q$ such that $q\equiv 1\mod M$, where $M=4N=2^{r+2}3^{r-1}$. The value of $M$ is consequence of our choice of size $N$ for the $\mathsf{DCT}$ earlier in Section 4.
In both cases, for all indices $i$ and $j$, we use
$2\cos\left(\frac{2\pi(2j+1)i}{4N}\right)=\zeta_{M}^{(2j+1)i}+\zeta_{M}^{-(2j+1)i}\in\mathbb{F}_{q}.$
@@
In fact, our multiplication algorithm extends naturally to a composite modulus $pq$ for two primes $p$ and $q$ satisfying the equivalence conditions above. This follows as a direct application of the Chinese Remainder Theorem. We use the ring isomorphism $\mathbb{Z}/pq\mathbb{Z}\cong\mathbb{F}_{p}\times\mathbb{F}_{q}$ to construct the required root of unity in $\mathbb{Z}/pq\mathbb{Z}$ that maintains its exact order under the projections to $\mathbb{F}_{p}$ and $\mathbb{F}_{q}$. This is enough to define the $\mathsf{DCT}$ and $\mathsf{IDCT}$ over $\mathbb{Z}/pq\mathbb{Z}$. Note that the existence of the inverse of $2$ and $N$ is also guaranteed by the isomorphism.
@@ 5 PLWE in $\mathbb{Z}[x]/(\Psi_{n}(x))$ via fast base change
Before we can prove Lemma 5.1, we need a few more results regarding the complexity of polynomial evaluation on the Chebyshev nodes $x_{j}$ (5) and Chebyshev interpolation. In [37, 38], the authors describe an algorithm to evaluate a polynomial $p(x)$ of degree at most $N-1$ at the Chebyshev nodes $x_{j}$ with complexity $\mathcal{O}(N\log N)$. In [37], this result is stated as the following theorem.
@@ 6 Small roots of cyclotomic polynomials modulo a prime
Let $\mathcal{O}_{f}=\mathbb{Z}[x]/(f(x))$ and $R_{q}=\mathcal{O}_{f}/q\mathcal{O}_{f}$ as before. The root-based attacks from [15] show that if the polynomial $f(x)$ has a small root modulo the reducing prime $q$, then there exists a polynomial time evaluation attack against the decisional PLWE problem. In this section, we compare heuristically cyclotomic fields and their maximal real subfields for attacks against roots of small residue. The numerical computations are based on the paper [21], where the PLWE–RLWE equivalence is proved for conductors of the form $4p$, where $p$ is a prime. For a reasonable range of parameters, our goal is to show that on average the minimal polynomials of the maximal real extensions are not worse than cyclotomic polynomials when it comes to having small roots. In practice, this means that we can easily find a pair of primes $(p,q)$ such that the polynomial $\Psi_{4p}(x)$ with prime modulus $q$ is not susceptible to attacks against roots of small residue.
We show that between the two families of minimal polynomials, $\Phi_{p}(x)$ and $\Psi_{4p}(x)$ of the same degree, the maximal real extensions $\Psi_{4p}(x)$ tend to have fewer small roots $x\in S=\{\pm 2,\pm 3,\pm 4,\pm 8\}$ when $p$ is large enough. On the other hand, for cyclotomic polynomials of small degree $p<200$ and reducing primes $q<5\cdot 10^{10}$, we did not encounter any small roots in $S$.
The roots in the set $S$ are small powers of two. We also included another small prime 3 because our preliminary computations suggested that it might occur as a root relatively often. This choice allowed us to study the frequency of the roots in $S$ to see if it varies a lot between the elements of $S$.
@@
  • $q=3329$ and $n=256$ for ML-KEM, formerly known as Kyber [1].
  • $q=8380417$ and $n=256$ for ML-DSA, formerly known as Dilithium [2].
  • $q=12289$ and $n=512$ in security level 1 and $n=1024$ in security level of Falcon [39].
@@
In practice, for PLWE instances the sparsity of the vulnerable polynomials allows us to pre-emptively test for small zeros for different prime moduli $q$ and choose a safe one. In this regard, the weak polynomials we present are only of theoretical interest in the PQC realm. However, it is worth mentioning that in [16], the authors attack the PLWE problem by using as the modulus the NTRU polynomials with parameters:
@@
The situation is slightly different in the homomorphic encryption setting, where much larger values of both the degree $n$ and the prime modulus $q$ can occur. For instance, within HELib [40], a large initial ciphertext modulus of 780 bits is used, which needs a cyclotomic degree of 55831. Thus, the ring is of dimension $\phi(55831)=54000$ achieving 129 bits of security. More specifically, in [41, Table 4.2] we can observe that $n$ runs in powers of 2 from 1024 to 131072, and $q$, depending on $n$, runs between 29 to 1920 bits. These parameters refer to power-of-two cyclotomic polynomials.
@@
For primes from $p<1500$, we computed the evaluations of $\Phi_{p}(x)$ and $\Psi_{4p}(x)$ at the elements of the set $S$ and then reduced the evaluations with a varying prime modulus $q$ running as high as $5\cdot 10^{10}$. Note that thank to the result from [20, 21], we could exclude the possibility of $\alpha=\pm 2$ being a root of $\Psi_{4p}(x)$ modulo any prime $q$. Furthermore, the minimal polynomials $\Psi_{4p}(x)$ are even [36] so $\alpha\in S$ is root if and only if $-\alpha\in S$ is a root. Thus, in our numerical search the number of roots at $\alpha$ and $-\alpha$ are the same for $\Psi_{4p}(x)$. Finally, note that for a fixed $p$, the degrees of the two polynomials are equal
@@
The counts for the number of zeros we found for the evaluation points in $S$ is given in Table 1. We did not observe a significant difference between the number of zeros at the small powers of 2 or at the small prime $3$. In Table 1 the number of zeros of $\Psi_{4p}(x)$ at $\alpha$ and $-\alpha$ are equal due to the even parity of the polynomials.